OpenStudy (anonymous):

The equation x"squared" + y"squared" - 2x + 2y -1 = 0 is the general form of the equation of a circle. What is the standard form of the equation?

OpenStudy (campbell_st):

complete the square in x and y \[x^2 -2x + y^2 + 2x = -1\] what needs to be added to make a perfect square in x and a prefect square in y. add those values to both sides of the equation then factor and the equation will be in the form \[(x - a)^2 + (y + b)^2 = r^2\] hope it helps

OpenStudy (anonymous):

can you explain it a lil more plzz